# Multi-Layered Reasoning from a Single Viewpoint for Learning See-Through Grasping

### v4

@@ Abstract
Sensory substitution enables biological systems to perceive stimuli typically obtained by another organ, which is inspirational for physical agents. Multi-modal perception of intrinsic and extrinsic interactions is critical in building an intelligent robot that learns. This study presents a Vision-based See-Through Perception (VBSeeThruP) architecture that simultaneously perceives multiple intrinsic and extrinsic modalities via a single visual input in a markerless way, all packed within a soft robotic finger using the Soft Polyhedral Network design. It is generally applicable to miniature vision systems placed underneath deformable networks with a see-through design, capturing real-time images of the network’s physical interactions induced by contact-based events overlayed on top of the visual scene of the external environment, as demonstrated in the ablation study. We present the VBSeeThruP’s capability for learning reactive grasping without using external cameras or dedicated force and torque sensors on the fingertips. Using the inpainted scene and the deformation mask, we further demonstrate the multi-modal performance of the VBSeeThruP architecture to simultaneously achieve various perceptions, including but not limited to scene inpainting, object detection, depth sensing, scene segmentation, masked deformation tracking, 6D force/torque sensing, and contact event detection, all within a single sensory input from the in-finger vision markerlessly.
Keywords Multi-Modal Learning Latent Representation Tactile Robotics
@@ 1 Introduction
Sensory substitution [1] is a technique where one sensory modality is used to provide information typically perceived by another. Biological systems utilize visual cues to infer contact-based information, serving as inspiration for achieving multi-modal robotic perception [2]. In classical settings, such as the “Rubber Hand Experiment” [3], research has shown that humans can translate contact-based sensations into visual or auditory signals, enabling them to see through touch [4] or hear textures [5]. This concept naturally extends to understanding deformable interactions in both human and biological systems, where visual observations of deformations can infer underlying contact-based interactions and material properties [6]. This principle forms the basis of Vision-Based Deformable Perception (VBDeformP), particularly in robotic manipulation, where the growing adoption of miniature vision sensors [7, 8] and robot learning methods [9, 10] are also notable. Accurately perceiving contact-based information is crucial for physical agents, such as robots, to effectively grasp and manipulate objects [11]. Leveraging visual cues of deformations offers a promising solution [12].
@@
Robotics research has increasingly explored vision-based techniques to understand deformable objects [13] and their interactions with the object-centric environment [14]. Simultaneously achieving a comprehensive perception of the scene (i.e., object detection, recognition, and spatial understanding) and the touch-based interactions (i.e., surface texture, material properties, contact forces) during object manipulation with robots is an active research area [10].
@@
However, achieving multi-modal perception within a single, compact visual input, especially when the sensor is embedded inside a soft robotic fingertip, poses a significant challenge [15]. The deformable interface between a miniature camera and the external environment plays a critical role in the accuracy and applicability of the system. Its design directly impacts the quality of both visual and contact-based information acquired, making it an emerging research area in tactile robotics [16]. An ideal system would be able to decouple the contact-based interaction gathered from the soft interface’s deformation from the miniature vision that also perceives the scene, all from a single visual input.
@@
This work presents a Vision-based See-Through Perception (VBSeeThruP) architecture that enables multi-layered reasoning to learn reactive grasping from a single visual input behind a deformable network, serving as a markerless representation of fingertips, as shown in Fig. 1. The VBSeeThruP approach addresses the challenges of concurrently perceiving both finger deformation induced by contact-based interactions and rigid objects within the scene, all within a unified visual input for physical agents. For enhanced interactions with omni-directional adaptation of objects, we adopted the Soft Polyhedral Network (SPN) reported in recent research [17, 18], and further achieved markerless representations through multi-modal perception. The miniature vision captures real-time images of the SPN’s physical deformation during contact-based interaction while preserving partial information of the external scene. Before contact, a static SPN mask is used to inpaint the occluded scene via a brief “head-up” camera movement. Subsequently, video object segmentation of the SPN’s large-scale deformations provides markerless contact-based perception by estimating 6D forces and torques during contact. We demonstrate the viability of this approach through an ablation study and a reactive grasping task, further showcasing its potential to perform object detection, depth sensing, and scene segmentation once occluded regions are inpainted. This offers a unified solution that could streamline traditional hand-eye systems into a single in-finger vision sensor.
@@ 2 Related Works
Vision underpins robotic perception by enabling object detection, localization, and scene understanding [19, 20]. However, properties such as stiffness, friction, and mass are difficult to infer from vision alone [21, 22], calling for integration of alternative modalities.
@@ 2.1 Differentiating Vision-based Sensing and Perception
| Example (Visual) | Camera pixels | Object detection, scene understanding |
| Example (Tactile) | Force/pressure readings | Recognizing object stiffness or shape |
Sensing provides the foundational information required by robots or intelligent agents to interact with the physical world. It refers to directly acquiring raw data from the environment using sensors. It is the preliminary stage of gathering signals or measurements. It typically involves hardware (e.g., cameras, microphones, force sensors, temperature sensors); produces unprocessed, quantitative, often noisy data; and measures passive or active physical signals without interpreting their meaning. For example, a camera captures pixel-level images, or a microphone records sound waves, which are typically considered forms of sensing. Perception, on the other hand, builds on sensing data, enabling robots to make informed decisions, adapt to new situations, and interact intelligently with the environment or humans. It involves processing, interpreting, and understanding sensory data to derive meaningful information or actionable insights about the environment or state. It often refers to the computational or cognitive processes involved in analyzing and interpreting sensor data, which include filtering, classification, recognition, inference, and decision-making tasks. Perception bridges the gap between raw sensing and actionable knowledge or behavior. For example, identifying objects within an image, recognizing human speech from audio signals, or estimating forces from tactile sensor data are generally considered forms of perception.
@@ 2.2 A General Classification of Vision-based Perception
An object-centric representation [23] leads to rigid-body assumptions or deformable interactions for visual understanding and reasoning, promoting robotic vision via Vision-based Rigid Perception (VBRigidP) and Vision-based Deformable Perception (VBDeformP), which are the leading approaches.
@@ 2.2.1 Vision-based Rigid Perception (VBRigidP)
VBRigidP methods traditionally assume rigid objects, leveraging tools such as Faster R-CNN [24] or YOLO [25] for detection, combined with pose estimation techniques (PnP, ICP) [26, 27] for manipulation tasks [28]. Although effective in structured industrial settings, VBRigidP and the rigid-object assumption become limiting for real-world scenarios involving deformable or partially deformable objects.
@@ 2.2.2 Vision-based Deformable Perception (VBDeformP)
VBDeformP extends beyond rigid bodies by tracking non-rigid shapes [29, 30]. Soft robotics, cloth manipulation, and surgical robotics applications exploit visual deformation cues to infer forces or material properties [31, 32]. By analyzing deformation in video, robots can estimate contact forces [33], plan tasks for large-scale cloth handling [31], or reconstruct object geometry [34], all without the need for specialized tactile sensors.
Emerging from VBDeformP is Vision-based Tactile Perception, which interprets tactile interaction by analyzing deformation images [35, 21]. Research has shown that detailed force and contact geometry can be inferred from images of a soft interface undergoing compression [36, 35], which can be further embedded into multi-fingered systems for object manipulation [37, 38]. Recent research has demonstrated tactile perception inferred at the sensor (e.g., normal and tangential forces [35], vibration [39], thermal [21], pretouch proximity [40]), contact (e.g., contact geometry [35], force and torque [18], contact events [41], material properties [42]), object (e.g., object localization [43], shape [33, 44], mass and dynamics [36], contents of containers [45]), and action (e.g., can be inferred from previous perceptual modalities for action selection and initialization, tactile feedback for low-level control, action termination, action outcome detection, and action outcome verification) levels [46], providing a cost-effective solution for tactile robotics [47].
It should be noted that the Vision-based Tactile Perception mentioned above is typically referred to as Vision-based Tactile Sensing (VBTS) in the computer vision and robotics literature. However, such terminology has caused some confusion among researchers in the cognitive science field, as the raw output signals obtained are camera pixels, and the subsequent tactile information is computationally processed to infer contact-based meanings for object manipulation. In addition, it could be argued that the tactile data collected using the VBTS method is not directly measured from compression readings regarding force or pressure on the interaction surface, but rather processed algorithmically based on the raw image pixels, making it more of a perception method than a sensing method. Without losing generality, we describe this approach as a subcategory of the VBDeformP method in this work, following the general review elaborated in Sec. 2.1.
@@ 2.2.3 Vision-based See-Through Perception (VBSeeThruP)
In this work, we introduce Vision-based See-Through Perception (VBSeeThruP), a novel approach that combines the benefits of both VBRigidP and VBDeformP to achieve multi-modal perception from a single visual input by reasoning about information obtained from different image layers during in-finger imaging. Cognitive science research on sensory substitution inspired our research in VBSeeThruP, as we can use a single visual sensing to infer multiple perceptual modalities by reasoning between various layers of information the camera captures. For example, our proposed VBSeeThruP architecture involves inferring contact-related information by tracking masked deformation from the See-Thru-Network (which is not limited to the Soft Polyhedral Network design exemplified in this work). The VBSeeThruP can also infer visual-related information after scene inpainting and the use of foundational models for a rich set of perceptual modalities in an object-centric scene. The combined effect is not achievable using VBRigidP or VBDeformP methods alone, and the resultant effect is enabled by introducing a See-Thru-Network in front of the in-finger camera. For example, when the network structure is 100% filled without see-through capabilities, it becomes equivalent to a VBTactileP (or VBTS) system. When it is 0% filled with complete see-through capabilities, it becomes equivalent to a VBRigidP system.
2.3 Multi-Modal Perception in Robotics with Vision
Combining vision and touch provides richer representations, improving classification [48], pose estimation [49], and in-hand manipulation [50]. This fusion is instrumental in cluttered or occluded environments [51, 52]. Pairing visual and auditory cues boosts performance in human action classification and material recognition [53, 54]. Multi-sensory attention mechanisms (vision, audio, touch) enhance performance for complex activities, such as pouring or packing [55, 56]. Recent multi-modal large language models (LLMs) combine visual perception with language-based reasoning, enabling zero-shot and few-shot policy learning [57, 58]. These systems exhibit stronger physical reasoning when additional modalities are incorporated [59, 60], widening opportunities for flexible, data-driven manipulation [61].
@@ 3 Proposed Methods
Our approach begins by denoting an image captured by a camera at timestep . The and are the user-defined output settings for the image size, measured in the total count of pixel rows and columns. Depending on the camera selection, the th row and th column pixel entry of the image may contain the brightness intensity of a single channel for grayscale images () or three channels in red, green, and blue (RGB) for color images (). Each pixel entry usually stores information in an 8-bit format, offering 256 conditions ranging from 0 to 255. Furthermore, we can add another data channel () to the RGB image to include the transparency state of each pixel.
@@ 3.1 Markerless Design of the See-Thru-Network
Figure 2 is the Hand-Eye system adopted in this study, built with aluminum extrusion based on the DeepClaw workstation [62], housing a collaborative robot (UR10e from Universal Robots) on a pedestal and a tabletop in front covered by black fabrics.
Although an “Eye” camera is fixed on the post shown in Fig. 2A, it is intentionally unplugged (therefore with a strikethrough) as we intend to implement visual perception of the scene using in-finger vision only behind a soft fingertip featuring a See-Through design using deformable materials. Figure 2B shows an enlarged view of the “Hand,” a two-fingered gripper (Model AG-160-95 from DH-Robotics) with its rigid fingertips replaced by the soft ones. Each soft fingertip contains a soft, omni-adaptive, markerless structure based on a variant of the Soft Polyhedral Network designs [63], mounted on a camera housing 3D-printed by UV-curable transparent resin (Somos WaterShed XC1112), where a miniature camera (Chengyue WX605 from Weixinshijie) is fixed inside.
As shown in Fig. 2C, the base plate can be customized to accommodate the installation of soft fingertips on force/torque (FT) sensors for testing or to use a gripper as a fingertip for soft and adaptive grasping. Figure 2D shows the objects used for testing in this study, including a Pringles Can from the Yale-Columbia-Berkley (YCB) object sets [64] and others commonly available at research labs, including a spray can, mineral water bottles filled halfway with water, a medicine bottle, a Bosche electrical screwdriver, pouch cell. Although both soft fingertips are functional, we only used the one whose viewing range partially covers the table at every frame throughout the following experiments.
3.2 Formalizing Multi-Modal VBSeeThruP

### v5

@@ Abstract
Sensory substitution enables biological systems to perceive stimuli that are typically perceived by another organ, which is inspirational for physical agents. Multimodal perception of intrinsic and extrinsic interactions is critical in building an intelligent robot that learns. This study presents a Vision-based See-Through Perception (VBSeeThruP) architecture that simultaneously perceives multiple intrinsic and extrinsic modalities from a single visual input, in a markerless manner, all packed into a soft robotic finger using the Soft Polyhedral Network design. It is generally applicable to miniature vision systems placed beneath deformable networks with a see-through design, capturing real-time images of the network’s physical interactions induced by contact-based events, overlaid on the visual scene of the external environment, as demonstrated in the ablation study. We present the VBSeeThruP’s capability for learning reactive grasping without using external cameras or dedicated force and torque sensors on the fingertips. Using the inpainted scene and the deformation mask, we further demonstrate the multimodal performance of the VBSeeThruP architecture to simultaneously achieve various perceptions, including but not limited to scene inpainting, object detection, depth sensing, scene segmentation, masked deformation tracking, 6D force/torque sensing, and contact event detection, all within a single sensory input from the in-finger vision markerlessly.
keywords:
Multimodal Learning , Latent Representation , Tactile Robotics[SUSTech]organization=Southern University of Science and Technology,city=Shenzhen, postcode=518055, state=Guangdong, country=China
[MBZUAI]organization=Mohamed bin Zayed University of Artificial Intelligence,country=United Arab Emirates
@@ 1 Introduction
Sensory substitution (Bach-y-Rita and W. Kercel, 2003) is a technique where one sensory modality is used to provide information typically perceived by another. Biological systems utilize visual cues to infer contact-based information, providing inspiration for achieving multimodal robotic perception (Caspar et al., 2015). In classical settings, such as the “Rubber Hand Experiment” (Botvinick and Cohen, 1998), research has shown that humans can translate contact-based sensations into visual or auditory signals, enabling them to see through touch (Tsakiris and Haggard, 2005) or hear textures (Lundborg et al., 1999). This concept naturally extends to understanding deformable interactions in both human and biological systems, where visual observations of deformations can infer underlying contact-based interactions and material properties (Yin et al., 2021). This principle forms the basis of Vision-Based Deformable Perception (VBDeformP), particularly in robotic manipulation, where the growing adoption of miniature vision sensors (Li et al., 2018; Alspach et al., 2019) and robot learning methods (Lloyd and Lepora, 2024; Bauza et al., 2024) are also notable. Accurately perceiving contact-based information is crucial for physical agents, such as robots, to effectively grasp and manipulate objects (Mittendorfer and Cheng, 2011). Leveraging visual cues of deformations offers a promising solution (Lambeta et al., 2024).
@@
Robotics research has increasingly explored vision-based techniques to understand deformable objects (Weng et al., 2024) and their interactions with the object-centric environment (Zhu et al., 2021). Simultaneously achieving a comprehensive perception of the scene (i.e., object detection, recognition, and spatial understanding) and the touch-based interactions (i.e., surface texture, material properties, contact forces) during object manipulation with robots is an active research area (Bauza et al., 2024).
@@
However, achieving multimodal perception from a single, compact visual input, especially when the sensor is embedded in a soft robotic fingertip, poses a significant challenge (Fazeli et al., 2019). The deformable interface between a miniature camera and the external environment is critical to the system’s accuracy and applicability. Its design directly impacts the quality of both visual and contact-based information acquired, making it an emerging research area in tactile robotics (Sun et al., 2022). An ideal system would be able to decouple the contact-based interaction inferred from the soft interface’s deformation from the miniature vision that also perceives the scene, using a single visual input.
@@
This work presents a Vision-based See-Through Perception (VBSeeThruP) architecture that enables multi-layered reasoning to learn reactive grasping from a single visual input behind a deformable network, serving as a markerless representation of fingertips, as shown in Fig. 1. The VBSeeThruP approach addresses the challenges of concurrently perceiving both finger deformation induced by contact-based interactions and rigid objects within the scene, all within a unified visual input for physical agents. For enhanced interactions with omni-directional adaptation of objects, we adopted the Soft Polyhedral Network (SPN) reported in recent research (Guo et al., 2024a; Liu et al., 2024b), and further achieved markerless representations through multimodal perception. The miniature vision captures real-time images of the SPN’s physical deformation during contact-based interaction while preserving partial information of the external scene. Before contact, a static SPN mask is used to inpaint the occluded scene via a brief “head-up” camera movement. Subsequently, video object segmentation of the SPN’s large-scale deformations provides markerless contact-based perception by estimating 6D forces and torques during contact. We demonstrate the viability of this approach through an ablation study and a reactive grasping task, further demonstrating its potential for object detection, depth sensing, and scene segmentation once occluded regions are inpainted. This offers a unified solution that could streamline traditional hand-eye systems into a single in-finger vision sensor.
@@ 2 Related Works
Vision underpins robotic perception by enabling object detection, localization, and scene understanding (Kirillov et al., 2023; Robinson et al., 2023). However, properties such as stiffness, friction, and mass are difficult to infer solely from vision (Li et al., 2024a; Zhao et al., 2025), calling for the integration of alternative modalities.
@@ 2.1 Differentiating Vision-based Sensing and Perception
Sensing provides the foundational information that robots and intelligent agents need to interact with the physical world. It refers to directly acquiring raw data from the environment using sensors. It is the preliminary stage of gathering signals or measurements. It typically involves hardware (e.g., cameras, microphones, force sensors, temperature sensors); produces unprocessed, quantitative, often noisy data; and measures passive or active physical signals without interpreting their meaning. For example, a camera captures pixel-level images, and a microphone records sound waves, both of which are typically considered forms of sensing. Perception, on the other hand, builds on sensing data, enabling robots to make informed decisions, adapt to new situations, and interact intelligently with the environment or humans. It involves processing, interpreting, and understanding sensory data to derive meaningful information or actionable insights about the environment or state. It often refers to the computational or cognitive processes involved in analyzing and interpreting sensor data, including filtering, classification, recognition, inference, and decision-making. Perception bridges the gap between raw sensing and actionable knowledge or behavior. For example, identifying objects within an image, recognizing human speech from audio signals, or estimating forces from tactile sensor data are generally considered forms of perception.
@@ 2.2 A General Classification of Vision-based Perception
An object-centric representation (Kroemer et al., 2021) leads to rigid-body assumptions or deformable interactions for visual understanding and reasoning, promoting robotic vision via Vision-based Rigid Perception (VBRigidP) and Vision-based Deformable Perception (VBDeformP), which are the leading approaches.
@@ 2.2.1 Vision-based Rigid Perception (VBRigidP)
VBRigidP methods traditionally assume rigid objects, leveraging tools such as Faster R-CNN (Ren et al., 2017) or YOLO (Redmon et al., 2016) for detection, combined with pose estimation techniques (PnP, ICP) (Lepetit et al., 2009; Besl and McKay, 1992) for manipulation tasks (Bohg et al., 2014). Although effective in structured industrial settings, VBRigidP and the rigid-object assumption become limiting for real-world scenarios involving deformable or partially deformable objects.
@@ 2.2.2 Vision-based Deformable Perception (VBDeformP)
VBDeformP extends beyond rigid bodies by tracking non-rigid shapes (Arriola-Rios et al., 2020; Kroemer et al., 2011). Soft robotics, cloth manipulation, and surgical robotics applications exploit visual deformation cues to infer forces or material properties (Zhu et al., 2022; Deng and Hsu, 2024). By analyzing deformation in video, robots can estimate contact forces (Luu et al., 2023), plan tasks for large-scale cloth handling (Zhu et al., 2022), or reconstruct object geometry (Guo et al., 2024b), all without specialized tactile sensors.
Emerging from VBDeformP is Vision-based Tactile Perception, which interprets tactile interaction by analyzing deformation images (Yuan et al., 2017a; Li et al., 2024a). Research has shown that detailed force and contact geometry can be inferred from images of a soft interface undergoing compression (Wang et al., 2022; Yuan et al., 2017a), which can be further embedded into multi-fingered systems for object manipulation (Liu and Adelson, 2022; Zhao and Adelson, 2023). Recent research has demonstrated tactile perception inferred at the sensor (e.g., normal and tangential forces (Yuan et al., 2017a), vibration (Kent et al., 2021), thermal (Li et al., 2024a), pretouch proximity (Hogan et al., 2022)), contact (e.g., contact geometry (Yuan et al., 2017a), force and torque (Liu et al., 2024b), contact events (Zhang et al., 2020), material properties (Yuan et al., 2017b)), object (e.g., object localization (Xu et al., 2024), shape (Luu et al., 2023; Van Duong and Ho, 2021), mass and dynamics (Wang et al., 2022), contents of containers (Zhang et al., 2021)), and action (e.g., can be inferred from previous perceptual modalities for action selection and initialization, tactile feedback for low-level control, action termination, action outcome detection, and action outcome verification) levels Li et al. (2020), providing a cost-effective solution for tactile robotics Haddadin et al. (2019).
It should be noted that the Vision-based Tactile Perception mentioned above is typically referred to as Vision-based Tactile Sensing (VBTS) in the computer vision and robotics literature. However, such terminology has caused some confusion among researchers in the cognitive science field, as the raw output signals obtained are camera pixels, and the subsequent tactile information is computationally processed to infer contact-based meanings for object manipulation. In addition, it could be argued that the tactile data collected using the VBTS method is not directly measured from compression readings of force or pressure on the interaction surface, but rather processed algorithmically from raw image pixels, making it more of a perception method than a sensing method. Without losing generality, we describe this approach as a subcategory of the VBDeformP method in this work, following the general review elaborated in Sec. 2.1.
@@ 2.2.3 Vision-based See-Through Perception (VBSeeThruP)
In this work, we introduce Vision-based See-Through Perception (VBSeeThruP), a novel approach that combines the benefits of both VBRigidP and VBDeformP to achieve multimodal perception from a single visual input by reasoning over information from different image layers during in-finger imaging. Cognitive science research on sensory substitution inspired our research in VBSeeThruP, as we can use a single visual sensor to infer multiple perceptual modalities by reasoning across the various layers of information the camera captures. For example, our proposed VBSeeThruP architecture infers contact-related information by tracking masked deformation in the See-Thru-Network (which is not limited to the Soft Polyhedral Network design exemplified in this work). The VBSeeThruP can also infer visual-related information after scene inpainting and the use of foundational models for a rich set of perceptual modalities in an object-centric scene. The combined effect is not achievable with VBRigidP or VBDeformP alone, and it is enabled by introducing a See-Thru-Network in front of the in-finger camera. For example, when the network is 100% filled and does not have see-through capabilities, it becomes equivalent to a VBTactileP (or VBTS) system. When it is 0% (or fully transparent), it becomes equivalent to a VBRigidP system.
2.3 Multimodal Perception in Robotics with Vision
Combining vision and touch provides richer representations, improving classification (Falco et al., 2017), pose estimation (Murali et al., 2024), and in-hand manipulation (Suresh et al., 2024). This fusion is instrumental in cluttered or occluded environments (Dutta et al., 2025; Navarro-Guerrero et al., 2023). Pairing visual and auditory cues boosts performance in human action classification and material recognition (Pieropan et al., 2014; Dimiccoli et al., 2022). Multi-sensory attention mechanisms (vision, audio, touch) enhance performance for complex activities, such as pouring or packing (Li et al., 2023; Wang et al., 2023). Recent multimodal large language models (LLMs) combine visual perception with language-based reasoning, enabling zero-shot and few-shot policy learning (Li et al., 2024b; Jiang et al., 2023). These systems exhibit stronger physical reasoning when additional modalities are incorporated (Lai et al., 2024; Driess et al., 2023), widening opportunities for flexible, data-driven manipulation (Liu et al., 2024a).
@@ 3 Proposed Methods
Our approach begins by denoting an image captured by a camera at timestep . The and are the user-defined output settings for the image size, measured in the total count of pixel rows and columns. Depending on the camera selection, the th row and th column pixel entry of the image may contain the brightness intensity of a single channel for grayscale images () or three channels in red, green, and blue (RGB) for color images (). Each pixel entry typically stores information in an 8-bit format, offering 256 levels from 0 to 255. Furthermore, we can add another data channel () to the RGB image to encode each pixel’s transparency state.
@@ 3.1 Markerless Design of the See-Thru-Network
Figure 2 is the Hand-Eye system adopted in this study, built with aluminum extrusion based on the DeepClaw workstation (Wan et al., 2020), housing a collaborative robot (UR10e from Universal Robots) on a pedestal and a tabletop in front covered by black fabrics.
Although an “Eye” camera is fixed on the post shown in Fig. 2A, it is intentionally unplugged (therefore with a strikethrough) as we intend to implement visual perception of the scene using in-finger vision only behind a soft fingertip featuring a See-Through design using deformable materials. Figure 2B shows an enlarged view of the “Hand,” a two-fingered gripper (Model AG-160-95 from DH-Robotics) with its rigid fingertips replaced by the soft ones. Each soft fingertip contains a soft, omni-adaptive, markerless structure based on a variant of the Soft Polyhedral Network designs (Wan et al., 2022), mounted on a camera housing 3D-printed by UV-curable transparent resin (Somos WaterShed XC1112), where a miniature camera (Chengyue WX605 from Weixinshijie) is fixed inside.
As shown in Fig. 2C, the base plate can be customized to accommodate the installation of soft fingertips on force/torque (FT) sensors for testing or to use a gripper as a fingertip for soft and adaptive grasping. Figure 2D shows the objects used for testing in this study, including a Pringles Can from the Yale-Columbia-Berkley (YCB) object sets (Calli et al., 2015) and others commonly available at research labs, including a spray can, mineral water bottles filled halfway with water, a medicine bottle, a Bosche electrical screwdriver, pouch cell. Although both soft fingertips are functional, we used only the one whose field of view partially covers the table at every frame in the following experiments.
3.2 Formalizing Multimodal VBSeeThruP